10.) A 32 mL solution of HNO3 has a concentration of 0.476 M. Calculate the pH and then calculate the mass in grams of HNO3 in solution.
1
Expert's answer
2020-05-19T08:44:00-0400

C=n/V=> n=C*V=0,476*0,032=0,015 moll

m(HNO3)=n*Mr= 0,015*63=0,945g

pH=-lg[H+] => HNO3= H+ + NO3-

=>C(H+)=C(HNO3)=>

pH= -lg(0,476)=0,322